Find the x and y intercepts and graph the equation plotting the intercepts. Please show all of your work. -3x-7y=21
Answer by mananth(16949)

-3 x + -7 y = 21
-7 y = 21 x+ 3 x
y= -3 -0.43 x
when x = 0 x= 0 ,( 0 , -3 )
When y=0 x= -7 ,( -7 , 0 )
When x= -7 , y= 0 ( -7 , 0
When x= 0 , y= 0 ( 0 , 0 )
When x= 7 , y= 0 ( 7 , 0 )